The perimeter of a rectangle is 47 cm more than the width of the rectangle. the length is 16 cm. what's the width of the rectangle?
Answer by mananth(16946)

let width be x
length =16
perimeter = x+47
...
2*(x+16)=x+47
2x+32=x+47
x=47-32
x=15 the width
...
m.ananth@hotmail.ca